You own a portfolio that has $2,100 invested in Stock A and $3,650 invested in Stock B. If the expected returns on these stocks are 8 percent and 15 percent, respectively, what is the expected return on the portfolio?
Answer #1

Total investment = investment in A + investment in B

Total investment = 2100 + 3650

Total investment = 5750

_________________________________________

Weight of A = 2100/5750 = 0.365217

Weight of B = 3650/5750 = 0.634783

________________________________________

Expected return = weight*return

Expected return = (8%*0.365217)+(15%*0.634783)

Expected return = 2.92% + 9.52%

Expected return = 12.44%